### 11 Mar 2024 v1

# Transparent AI Disclosure Obligations: Who, What, When, Where, Why, How

Abstract.
Advances in Generative Artificial Intelligence (AI) are resulting in AI-generated media output that is (nearly) indistinguishable from human-created content. This can drastically impact users and the media sector, especially given global risks of misinformation. While the currently discussed European AI Act aims at addressing these risks through Article 52’s AI transparency obligations, its interpretation and implications remain unclear. In this early work, we adopt a participatory AI approach to derive key questions based on Article 52’s disclosure obligations. We ran two workshops with researchers, designers, and engineers across disciplines (N=16), where participants deconstructed Article 52’s relevant clauses using the 5W1H framework. We contribute a set of 149 questions clustered into five themes and 18 sub-themes. We believe these can not only help inform future legal developments and interpretations of Article 52, but also provide a starting point for Human-Computer Interaction research to (re-)examine disclosure transparency from a human-centered AI lens.
EU AI Act, Article 52, generative artificial intelligence, disclosures, transparency, questions
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
“Fake realities will create fake humans. Or, fake humans will generate fake realities
and then sell them to other humans, turning them, eventually, into forgeries of themselves.”
— Philip K. Dick, I Hope I Shall Arrive Soon, 1980
The field of Artificial Intelligence (AI), and the media sector at large, are undergoing a transformative change. Foundation models, developed on the basis of deep neural networks and self-supervised learning, have gained widespread acceptance (Bommasani et al., 2021). These models have led to the emergence of Generative AI (GenAI) tools like Midjourney and ChatGPT, and have shown impressive proficiency in producing media content, including images (Saharia et al., 2022), text (OpenAI, 2023), videos (Ho et al., 2022), and audio (Borsos et al., 2022). These have the potential to drastically impact users and the media sector, essentially blurring the line between fiction and reality as users engage with media. Model output has reached a level by which humans can no longer perceive GenAI outputs as distinguishable from human-generated content (Groh et al., 2023). What digital media appears to be true and authentic cannot necessarily be trusted, and is reportedly fueling the spread of mis- and disinformation [^1] https://www.axios.com/2023/02/21/chatbots-misinformation-nightmare-chatgpt-ai. Therefore it is imperative to disclose the use of AI in the generation/manipulation of media content. According to the World Economic Forum’s 2024 Global Risks Report (2024, 2024), “misinformation and disinformation” were ranked as the highest global risk anticipated over the next two years. This also comes at a time where OpenAI issues a statement regarding the upcoming 2024 worldwide elections (OpenAI, 2024), emphasizing: “Transparency around AI-generated content”. In fact, in the European Union (EU), such an obligation will be incorporated into the upcoming AI Act (Act, 2022). Despite that the AI Act is still in development and subject to change, the core obligation to “disclose AI-driven interaction” appears to be a stable one.
Algorithmic approaches to automatically detect text (co-)produced using GenAI tools (e.g., ChatGPT) are so far unreliable (Sadasivan et al., 2023), and can result in consistent biases against specific user groups (e.g., non-native English writers (Liang et al., 2023)). While measures to encode provenance cryptographically in images (cf., Coalition for Content Provenance and Authenticity (C2PA, 2024)) or audio (SynthID (DeepMind, 2024)) are being taken, these are not widely implemented. Moreover, it remains unclear how these need to be displayed to users. Despite concerted efforts worldwide, constructing even simple disclosure measures (e.g., watermarks) that actually succeed in reducing public risks is itself a difficult task. Disclosure is defined as “to make known or public” (Webster, 2024), where in the context of AI, disclosures can “contain information about the data collection, data processing, and decision-making practices of a digital product and are voluntarily provided by the product’s vendor (an individual developer or an organization).” (Open Ethics Initiative (Lukianets, [n.d.])). A fundamental challenge here is in ensuring effective transparency measures, given the rapid pace of AI development and deployment. To mitigate harms and risks, regulatory efforts such as the AI Disclosure Act of 2023 in the United States (Torres, 2023) and the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) aim to tackle this. In this work, we focus on the EU AI Act proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses the issue of AI system transparency [^2] Even though the final text of the AI Act must still be agreed upon, the inclusion of a ”disclosure obligation” is (almost) a certitude..
In this early work, we adopt a human-centered approach to derive key questions that arise around transparent AI disclosures. We ask: RQ: What are the key considerations and concerns surrounding transparent AI disclosures in the context of the EU AI Act? We ran two participatory AI workshops with researchers, designers, and engineers across disciplines (N=16), where we utilized the 5W1H framework (Hart, 1996; Jia et al., 2016) to deconstruct the relevant clauses in Article 52 concerning AI disclosures for users and providers. This was done in the context of the media sector, with a focus on media consumption and production. Even if the language of Article 52 changes, our goal is to trigger reflection on disclosures and what these might look like. We contribute a set of 149 questions clustered into five themes and 18 sub-themes, that we anticipate can help drive interdisciplinary research forward in responsible AI. Our work aims at tackling the challenge of interpreting and implementing obligatory transparent AI disclosures for ever-evolving AI technology in an interdisciplinary manner. This not only helps inform legal developments and future interpretations of Article 52, but also provides a starting point for the Human-Computer Interaction (HCI) community to (re-)examine disclosures from a human-centered AI lens.
2. Background and Related Work
2.1. EU AI Act’s Article 52 and AI disclosure implications
We focus on the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses AI system transparency. It raises two important clauses:
[…]
From these clauses, the types of disclosure are unspecified, and terms such as ‘authentic’ are undefined, leaving them open for interpretation. For example, ‘what’, ‘how’, or ‘when’ should such AI disclosures take place, even if automatically detected, remains unclear. While the most recent draft of the AI Act (Monitor, [n.d.]) touches on aspects of this, where Art. 52(3) (second subparagraph) highlights when to disclose, much remains uncharted regarding human oversight processes when exceptions are made. We believe it is important to carefully assess the AI Act’s disclosure implications, before proceeding to Human-AI interactions, wherein explainability (Panigutti et al., 2023) may be an additional challenge. Furthermore, authenticity is itself a multidimensional concept (Kernis and Goldman, 2006), which raises the question of what aspects of inauthentic media need to be disclosed. AI disclosures, be they automatic (algorithmic) or provided by system designers or other stakeholders, may include the following: revealing when content is AI-generated, registering these emerging AI systems with a database, summarizing copyrighted material used in training these systems, publication of risk assessments, or even trust certification labels (Scharowski et al., 2023). Importantly, encounters with AI-generated content can impact the human experience of algorithms, and more broadly the psychology of Human-AI interaction (Sundar, 2020), to which we turn to next.
2.2. Human-AI interaction, media consumption, and transparent AI disclosures
For scholarly work, the Association for Computing Machinery has instated clear policies on GenAI, stating “The use of generative AI tools and technologies to create content is permitted but must be fully disclosed in the Work.” (ACM, 2024). This contributes to ongoing discussions on the ethics of AI disclosure in scholarly works (Hosseini et al., 2023). Within HCI research and practice, Schmidt et al. (Schmidt et al., 2024), in rethinking Human-Centered Design in the age of GenAI, emphasize the need for being “transparent and honest” when it comes to AI tool usage. Indeed, issues of ownership and agency arguably span the entire HCI research cycle (Elagroudy et al., 2024). Such interactions will continue to pervade not just scholarly discourse, but also the everyday media we consume, making it crucial to understand the impact on human perceptions. For algorithmic decision making, Langer et al. (Langer et al., 2022) showed that terminology (e.g., ‘algorithms’ vs. ‘artificial intelligence’) affects laypeople’s perceptions of system properties and evaluations (e.g., trust) – they recommend being mindful when choosing terms given unintended consequences, and their impact on HCI research robustness and replicability. Within COVID-19 health (mis-)information, Jia et al. (Jia et al., 2022) found that various misinformation labels (e.g., algorithm, community, third-party fact-checker) are dependent on people’s political ideology (liberal, conservative). Cloudy et al. (Cloudy et al., 2023) found that a news story presented as sourced from an AI journalist activated individuals’ machine heuristic (rule of thumb that machines are more secure and trustworthy than humans (Sundar and Kim, 2019)), which helps mitigate the hostile media bias effect. Furthermore, there may be hidden dangers in such labeling approaches, where they may lead users to believe that content that is not labeled is actually factual, when it may not be – the so-called “implied truth effect” (Pennycook et al., 2020). Epstein et al. (Epstein et al., 2023) found that participants consistently associated “AI Generated,” “Generated with an AI tool,” and “AI manipulated” with AI content out of nine potential labels, regardless of whether or not they were misleading. These works underscore the importance of disclosure-based interventions, and highlight the wild west of today’s disclosure approaches, from tool usage to perceiving and understanding disclosure labels to safeguard against fake news.
3. Method
3.1. Workshops: approach and objectives
To generate key questions in a participatory manner, we ran two workshops in December (2023) and January (2024). We utilized the 5W1H (Who, What, When, Where, Why, How) framework, which is widely used in journalism (Hart, 1996). Furthermore, this framework allows defining a high-dimensional design space (cf., software testing for cloud computing (Jia et al., 2016)), by which an initial set of research questions can be systematically formulated. Our objective was to help deconstruct the relevant clauses in Article 52 by identifying key questions related to the interpretation, implementation, and societal impact of AI disclosures for users and providers. This was done in the context of the media sector, with a focus on media consumption and production. We anticipated that this would help unravel core values, consideration, and risks behind each question type. Workshops were conducted at two different locations/institutes, each targeting a specific researcher demographic. At the first location (session 1), participants had expertise from law, political science, communication science, and artificial intelligence. At the second location (session 2), we had expertise from computer science, human-computer interaction, and (interaction) design. For each workshop session, we collected: demographics, informed consent, photos of the session, and the resulting questions generated.
3.2. Workshop materials
The workshop setup consisted of six flip charts that were attached to the walls of the workshop area (see Figure 1(b)). Each flip chart was given a heading of one of the 5W1H terms to indicate to participants which type of questions to ask there. Post-its and markers were provided. Article 52 and Article 3 were printed on an A4 landscape paper, and placed near each flip chart for quick reference. Article 3 provides definitions of ‘provider’ and ‘user’ [^3] In the latest draft version of the EU AI Act (leaked online on January 21, 2024) (Monitor, [n.d.]), ‘user’ is now referred to as ‘deployer’. However, this does not impact our workshop outcomes, even if participants interpreted ‘user’ as intended by the law (‘deployer‘) or as ‘recipient’., that helped participants have a clear understanding of what these ‘legal’ terms mean. Article 3 (current revision; bold added for emphasis) states:
[…]
3.3. Study procedure
Sessions began with a short  10 minute introduction to set the workshop context, to obtain participant consent (data protection and privacy), and for them to fill in a demographics form. Each workshop session consisted of three tasks, each lasting around 25-30 min., amounting to approximately one and a half hours in total. In the introduction, we first presented the EU AI Act, along with Article 52 and a brief explanation of disclosures. Participants were assigned an ID to note down with each post-it note. Participants were split into groups of two or three, depending on the total size in the session. The grouping was only for discussion and logistical purposes, so each question posed was still done individually. Task 1 and 2 were question generation tasks, where participants were instructed to individually generate questions for each of the 5W1H questions. After 10 minutes, they moved to a different question, such that at the end of  60 minutes each group has had the chance to provide their input for all six question types. Participants were also instructed to review the questions that were placed by participants on the flip chart, and to optionally vote with a +/- 1 post-it on the question if they were in agreement/disagreement. For Task 3, we divided each flip chart into two parts: users and providers. This provided participants the opportunity to reflect on and consider whether the posted questions were relevant for the user, the provider, or both (in which case it was placed in the center of the flip chart). Participants were offered candy in the session. At the end of each session, participants were thanked for their time and efforts.
3.4. Participants
We had 16 participants (8 female, 7 male, 1 non-binary), where the first workshop had seven, and the second nine participants. Twelve were in the 25-34 age group, three in the 18-24 age group, and one in the 35-44 age group. Three were pursuing their master’s degree, seven either their PhD or had research assistant roles, and six had completed their PhD with researcher roles (postdoctoral or higher). Participants’ affiliations were spread across five different institutes. Nine had expertise in computer science and engineering (incl. computer vision, natural language processing, and signal processing), two in HCI and design, two in communication science, two in law, and one in political science.
Figure 1. Snapshots of the workshop sessions showing participants generating questions.
(a) Workshop Session 1
(b) Workshop Session 2
Two images side by side showing participants in our workshops. The left image shows the workshop that took place at the first location targeting mainly those with a legal and policy background, and the second shows the workshop that look place at the second location, targeting mainly those with a computing and design background. In both photos, participants are writing on sticky notes on a flip-chart of a question type.
Figure 1. Snapshots of the workshop sessions showing participants generating questions.
3.5. Analysis approach
We analyzed our data using inductive thematic analysis (Clarke et al., 2015). First, the first author created early codes for the 155 questions. These codes were clustered into different topics (e.g., ‘misinformation’, ‘trust’, ‘ux’, ‘infovis’). Based on these, an initial set of sub-themes and themes emerged. The second author independently reviewed all questions and corresponding sub- and main themes. Instead of calculating statistical inter-rater reliability for the analysis, the consensus among the authors was reached through two online video discussions (each approximately 1.5 hours) (McDonald et al., 2019), where each question was carefully re-assessed by both researchers to arrive at a final list of sub-themes and themes.
4. Results
Our workshops resulted in a total of 155 questions, where the raw digitized data can be seen in Supplementary Material. After data cleaning (removing near-duplicates, irrelevant questions), this resulted in 149 questions. Questions were edited for typos, but left largely as is. Our thematic analysis resulted in five main themes, each with sub-themes. The full list of questions, themes, sub-themes, and participant votes are shown in Table 1 in the Appendix. We note that some questions may belong to more than one (sub-)theme, however for our analysis, we chose the most representative classification. Themes and sub-themes, with examples selected mainly based on participant votes, are presented below:
4.1. Theme 1: Ethical, Legal, and Policy Considerations
A primary theme (N=41) that emerged concerns the ethical, legal, and policy aspects of AI disclosures. This had three inter-related sub-themes. The first sub-theme covers the Ethical Implications of AI Use: “How much should you communicate? Does law provide any answers? If not, what is ethical?[Q1]; “How dangerous is the content generated by AI?[Q2]; “Why should this be a right?[Q3]. The second sub-theme consisted of questions pertaining to Legal Compliance and AI Disclosure: “What should be disclosed?[Q7]; “Who is going to be responsible for the consequences?[Q8]; “When should users be punished for not following the obligation? and when providers?[Q12]. The last sub-theme here consisted of questions related to Policy and Regulatory Impact. While some overlap may exist with legal compliance, questions under this sub-theme focused largely on the impact of such policies, rather than compliance per se. Examples here include: “Who is going to be impacted by non disclosure?[Q23]; “When do users need to disclose the use of AI? Consider continued influence effect[Q24]; “Who decides if a “context of use” is obvious enough to not require an “interaction”?[Q25]. These sub-themes show how disclosures (and their effectiveness) are dependent on a variety of inter-related factors, the resolution of which, too has ethical and legal implications (e.g., if disclosed, what should be disclosed and who should do so; responsibility being also legally defined, etc.). This has prompted recent work to discuss what GenAI regulation should focus on (cf., (Helberger and Diakopoulos, 2023; Hacker et al., 2023; Novelli et al., 2024)) and how such measures can be enforced – for example, given their multi-purpose usage, such general-purpose AI systems should consider safety from the onset, starting with data quality (Helberger and Diakopoulos, 2023).
4.2. Theme 2: Future Considerations, Evolving Context, & Practical Implementation
Another theme that emerged (N=19) concerns the future of AI technology and disclosures, their implementation, and their impact on society. This also had three sub-themes. The first sub-theme consisted of questions pertaining to the Evolving AI Technologies and Societal Impact, where key questions touched upon societal aspects where misinformation may be rampant, and included: “Who cares?[Q42]; “When can AI generated content affect real life affairs?[Q45]; “Where could fake AI content show up?[Q48]. The second sub-theme consisted of questions covering Future Trends and Legal Adaptation: “Why is it important to inform people?[Q49]; “Why should people care about this?[Q50]; “What is the definition of AI generated content? (e.g. images, text)[Q52]. The final sub-theme addresses questions related to Practical Challenges in AI Implementation: “How much effort should be put into the classification?[Q56]; “What kind of AI model could be used in such situations?[Q58]; “How to trigger user feedback for AI disclosure or interactions?[Q61]. Such considerations are important for better understanding the impact on society. For example, Yaqub et al. (Yaqub et al., 2020) confirmed that credibility indicators on social media can decrease the propensity to share fake news, however their impact varied, with fact checking services found to be most effective. Yet social media platforms come and go, with an evolving user base, which underscores the importance of accounting for dynamic (future) contexts.
4.3. Theme 3: Provider Responsibility and Industry Impact
Given that we focused on differences between users and providers, a theme (N=18) that emerged concerned the ethical responsibilities and obligations of the (media) provider. This consisted of three sub-themes. The first sub-theme concerns Ethical Considerations for Providers: “Why do providers need to disclose they use AI?[Q62]; “How should we treat users when not following the obligation? And how providers? (think punishment)[Q64]; “Why providers should consider disclosure in their decision making process?[Q68]. The second sub-theme covers Industry-Specific Impact and Challenges: “Why do providers want to use AI? (what is the benefit?)[Q68]; “What industry will be affected the most?[Q71]; “What is considered manipulation? Is there also a degree to this that is accounted for?[Q72]. The last sub-theme addresses Provider Disclosure Obligations: “Where to label information / content?[Q74]; “What freedom shall the provider have in formulating/offering information to the recipient?[Q77]; “What information should the provider disclose to the recipient of AI content?[Q78]. Given these, we highlight not just the ethical but also the business impact of AI disclosures, which may itself then impact ethical considerations. For example, in a field customer purchase setting using chatbots, Luo et al. (Luo et al., 2019) found that disclosure of chatbot identity before the machine–customer conversation took place reduced purchase rates by more than 79%, despite what is known about people’s machine heuristic (Sundar and Kim, 2019).
4.4. Theme 4: Trust, Authenticity, and User Empowerment
A key theme (N=33) of our workshops was on trust, authenticity, and how users can be empowered. This consisted of four sub-themes. The first sub-theme includes questions on Authenticity, Provenance, and Transparency: “How important is it for users to know the authenticity of media?[Q81]; “How are we going to verify authenticity?[Q82]; “Where can people manually verify if it is AI generated?[Q84]. The second sub-theme tackles the Human-centered AI issue of Building Trust in AI Systems: “How will different users interpret signals differently?[Q89]; “How does disclosure impact people’s trust and perceived credibility?[Q91]; “How to responsibly disclose artistic uses of AI?[Q92]. The third sub-theme focuses on Empowering Users through Education & Awareness: “How can we train users to spot authentic content without relying on explicit cues?[Q99]; “What type of information empowers/triggers “act of resistance”?[Q100]; “Where should we get ask if we suspect things (images, texts) are generated by AI?[Q103]. Finally, the last sub-theme focuses on User Responsibility and Agency: “Why do users need to know they are interacting with AI?[Q110]; “What information should the users acquire? Are they responsible for it?[Q111]; “Why is it important for promoting democracy?[Q112]. The sub-themes here touch upon important prior work that aims to address authenticity and trust, in the context of media. Scharowski et al. (Scharowski et al., 2023) examined the potential for AI certification labels (e.g., the ”Digital Trust Label” by the 2023 Swiss Digital Initiative), and found that these can mitigate data-related concerns expressed by end-users (e.g., privacy and data protection), however other concerns (e.g., model performance) remain challenging to address. Longoni et al. (Longoni et al., 2022) found that people rated news headlines written by AI as less accurate than those written by humans. Recently, Toff & Simon (Toff and Simon, 2023) found that on average audiences perceive news labeled as AI-generated as less trustworthy, even when the articles themselves are not evaluated as any less accurate or unfair. However, these effects largely stem from those whose pre-existing levels of trust in news are higher to begin with and among those knowledgeable about journalism.
4.5. Theme 5: User Experience, Information Overload, and Personalization
Our final theme (N=34) concerned all matters related to User Experience (UX), information overload, and personalization. The first sub-theme concerns questions on Personalization and User Preferences: “How much freedom should the users get to (not) see the classification?[Q115]; “How to adapt AI disclosures personally?[Q117]; “How to present them in different devices?[Q118]. The second sub-theme addresses questions on the Psychology of Human-AI Interaction: “Why do users trust AI more/less than a human?[Q120]; “How would the disclosure affect the users?[Q121]; “What is the psychological difference between AI and Human?[Q122]. The third sub-theme covers questions on Standardization: “What is reliable method to test the classification?[Q126]; “Where on the website should Al be disclosed? Terms of service, every piece of content, every element?[Q127]; “Who determines how disclosure should be provided, according to which standards?[Q128]. The fourth sub-theme focuses on User Interfaces and Information Overload: “Where is the classification going to be displayed?[Q133]; “Where is the balance between communicating enough and information overload?[Q134]; “How can you effectively communicate use of Al without distracting from content?[Q135]. The final sub-theme concerns User-Centric Information Design: “Who are you communicating to? e.g., different people may require different types of communication mechanisms[Q139]; “What (under-) information is meaningful for the recipient?[Q140]; “Where to provide the AI disclosure? Before the generated content or after that?[Q142]. Indeed, with respect to UX and personalization aspects of disclosure, the level of granularity by which such disclosures are shown (e.g., do they pertain to the training data or to surface interaction), any corresponding explanations, and the extent to which these increase trust, can further compound public risk perceptions. In this direction, in studying progressive disclosures for algorithmic decisions made by intelligent systems, Springer & Whittaker (Springer and Whittaker, 2020) found that users may benefit from initially simplified feedback that hides potential system errors, and thereafter assists users in building working system operation heuristics. This is line with Muraldihar et al.’s (Muralidhar et al., 2023) survey that found a simplified explanation of the AI system is sufficient for ensuring transparency, allow users to have a better sense of system accuracy, fairness, and privacy.
5. Discussion and Conclusion
5.1. Limitations and future work
First, AI technology is advancing rapidly, and our participant responses were likely geared toward their experiences with GenAI technology today. However in the future, more transparent and robust AI systems may afford different questions than what we observed. Relatedly, the law is not immutable – this is continuously evolving, and the AI Act may develop to cover some of the key questions raised in this work. Lastly, while we took care to ensure our question set and themes are exhaustive for the study of AI disclosures, there may be aspects that were missed. Similarly, themes may be decomposed further – however, for our current scope, we believe these themes have sufficient coverage and detail to inform future work. Lastly, these should not be seen as a final but rather organic, ever-growing list of questions to systematically guide research on transparent AI disclosures.
5.2. Transparent AI disclosures in and beyond the AI Act
Our work aimed at tackling the challenge of interpreting and implementing transparent AI disclosures for continuously evolving AI technology, through the lens of Article 52 as currently stated in the draft EU AI Act (Act, 2022). Through the 5W1H questions, we arrived at several themes and sub-themes – we believe each poses unique challenges for users, providers, and the media sector at large. With GenAI, media organizations are or will be undergoing a fundamental shift in practices. To echo Rakova et al. (Rakova et al., 2021), to better enable responsible AI work, organizations need to update their practices, which will require addressing both prevalent and emerging work practices. For transparent AI disclosures, as we observed from the diverse set of questions tackling multiple aspects of disclosures (from terminology to ethical and legal aspects to societal impact), we believe this is a complex, multi-faceted challenge (a ‘wicked’ problem) that requires concerted interdisciplinary efforts from researchers and practitioners across industries and institutes. As such, we believe on continually drawing on participatory AI and value-sensitive design approaches (Birhane et al., 2022; Ehsan and Riedl, 2020; Friedman, 1996) to creating AI for social good (Umbrello and Poel, 2021), ensuring human creative practices are safeguarded and can flourish (cf., (Li et al., 2024; Inie et al., 2023)). This means ensuring media organization practices and citizens (end users) are transparently informed when they interact with AI-generated media in a meaningful, usable manner, throughout the human-AI interaction timeline. As a step toward this change, we believe our contributed key questions drawn from wide-varying expertise can serve as a practical starting point for cross-discipline research in AI disclosure – by knowing what questions to ask, and for whom such questions may be most relevant, collaborative efforts can be strengthened.
Whether the most effective mechanism to tackle transparent AI disclosures constitutes watermarking to ensure copyright and traceable accountability (Tohline et al., 2008; Zhong et al., 2023), or the continued refinement of regulation and legal policy at an (inter-)national level (of which the AI Act is a strong pillar (Veale and Borgesius, 2021)), it is clear GenAI development needs to be acted upon responsibly - for and beyond media organization practice. For example, consider the use and misuse behind data privacy and security measures through interaction with consent banners (Gray et al., 2021), despite EU efforts such as in General Data Protection Regulation (GDPR) mandates. These can lend themselves to dark patterns and potentially malicious use of AI persuasion (Burtell and Woodside, 2023), raising serious ethical concerns pertaining to perceptual manipulations, cybersecurity, and the global risk of misinformation (cf., WEF’s 2024 Global Risks Report (2024, 2024)). To this end, we believe drawing on key human-centered principles (e.g., human well-being alignment, responsible design of transparent AI) of responsible AI development (Garibay et al., 2023) are of immediate necessity to protect citizens moving forward, and ensuring a well functioning democratic society.
5.3. Conclusion
Our work contributes 149 questions clustered into five themes and 18 sub-themes, that we believe can assist us in tackling the challenge of transparently communicating AI disclosures – for media providers, as well as end-users consuming media content everyday across platforms and devices. We hope our questions underscore the importance for better understanding user needs and reactions to transparency obligations, and support establishing user-centric designs for transparent AI disclosures that ultimately foster democratic societies based on truth rather than AI-generated fiction.
Acknowledgements.
We thank all our participants for participating in our workshops, and acknowledge their valuable contributions to this early work. This publication is part of the AI, Media & Democracy Lab (Dutch Research Council project number: NWA.1332.20.009).
Appendix A Appendix
Table 1. Themes, sub-themes, and participant vote counts for the set of questions generated during our workshops.
ID Questions Sub-theme Theme Votes
1
How much should you communicate? Does law provide any answers? If not what is ethical?
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
3
2
How dangerous is the content generated by AI?
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
2
3
Why should this be a right?
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
1
4
Why is it important to disclose the limitation and complexity of AI systems?
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
5
Whose interests have informed this law?
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
6
Why should similar obligation not apply to human in digital setting? (think “authenticity” in virtual worlds like 2nd life or WOW)”
Ethical Implications of AI Use
Ethical, Legal, and Policy Considerations
7
What should be disclosed?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
3
8
Who is going to be responsible for the consequences?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
3
9
Who is responsible for informing the user?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
3
10
Who is going to compensate for damages?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
1
11
Who decides what “interaction” means? Art 52.1 what does it mean?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
1
12
When should users be punished for not following the obligation? and when providers?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
1
13
When to update the law given the fast technological development?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
1
14
What are the intended uses and applications of AI systems as disclosed to users?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
15
When is disclosure not sufficient?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
16
When is communication authentic?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
17
Why we need to make laws for disclosing of it? Why we let the provider and the user to determine by themselves?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
18
Who should set the rule?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
19
Who has to go to prison if something happens?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
20
Who should contribute to develop the AI Disclosure?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
21
Who is responsible for AI Disclosure?
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
22
Who should be disclosed? Context
Legal Compliance and AI Disclosure
Ethical, Legal, and Policy Considerations
23
Who is going to be impacted by non disclosure?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
3
24
When do users need to disclose the use of AI? Consider continued influence effect
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
2
25
Who decides if a “context of use” is obvious enough to not require an “interaction”?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
2
26
What information should recipients not receive?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
2
27
When the AI disclosure should be introduced?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
2
28
What transparency level/degree needs to be provided to the users?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
1
29
Who is responsible for deciding if a disclosure is needed?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
1
30
When is it mandatory to disclose something is created by AI? Always?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
1
31
What is a good way to do AI disclosure in the context where the “mystery?” is needed?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
32
What is the content that AI should disclose?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
33
Who wants to keep informed about the fact that is generated by AI?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
34
Who is going develop/fine tune the specific AI to create such contents?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
35
Where in the value chain should transparency be provided?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
36
Where is this monitored from?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
37
Why differentiating AI and Human is necessary?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
38
When to introduce AI disclosure for different AI systems? Such as google search, chatGPT?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
39
What timeline is needed to adopt the new “disclosure” system?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
40
Who, within a chain of multiple actors, will be responsible for effectuating there disclosure obligations?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
41
When to provide AI disclosure for the users for different scenarios?
Policy and Regulatory Impact
Ethical, Legal, and Policy Considerations
42
Who cares?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
7
43
Who benefits from AI disclosure?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
2
44
When could fake AI content actually affect users’ opinion?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
45
When can AI generated content affect the real life affairs?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
46
Why did lawmakers opt for transparency/disclosures as the core solution to “inauthentic” content?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
47
When the AI disclosure agreement should work, avoid the effect of rapid development of AI technology?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
48
Where could fake AI content show up?
Evolving AI Technologies and Societal Impact
Future Considerations, Evolving Context, & Practical Implementation
49
Why is it important to inform people?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
2
50
Why should people care about this?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
2
51
Who is AI?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
2
52
What is the definition of AI generated content? (e.g. images, text)
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
53
Why researchers are interested in AI systems disclosure?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
54
Why would AI have access to real world data?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
55
What training data is required for AI to create fake content?
Future Trends and Legal Adaptation
Future Considerations, Evolving Context, & Practical Implementation
56
How much effort should be put into the classification?
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
2
57
Where the AI disclosure is not welcomed? (e.g., Academia, Industry)
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
58
What kind of AI model could be used in such situations?
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
59
How to disclose implicitly or explicitly?
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
60
How to correctly use AI for content production?
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
61
How to trigger user feedback for AI disclosure or interactions
Practical Challenges in AI Implementation
Future Considerations, Evolving Context, & Practical Implementation
62
Why do providers need to disclose they use AI?
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
3
63
Why wouldn’t the providers want to disclose that it’s an AI system?
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
2
64
How should we treat users when not following the obligation? And how providers? (think punishment)
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
1
65
Who determines the intention of the interaction?
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
66
What type of content should not be labeled?
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
67
Why provider should consider disclosure in their decision making process?
Ethical Considerations for Providers
Provider Responsibility and Industry Impact
68
Why do providers want to use AI? (what is the benefit?)
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
3
69
What is the effort required to disclose the information correctly?
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
2
70
When should providers start thinking about disclosure mechanisms in the production process?
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
71
What industry will be affected the most?
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
72
What is considered manipulation? Is there also a degree to this that is accounted for?
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
73
What exactly is a “deep fake”?
Industry-Specific Impact and Challenges
Provider Responsibility and Industry Impact
74
Where to label information / content?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
1
75
How can you know that the use of AI was not disclosed?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
1
76
When to not disclose information?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
1
77
What freedom shall the provider have in formulating/offering information to the recipient?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
1
78
What information should the provider disclose to the recipient of AI content
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
79
Why would media/press announce that they are using AI?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
80
How to decide to what extent the provider does AI disclosure?
Provider Disclosure Obligations
Provider Responsibility and Industry Impact
81
How important is it for users to know the authenticity of media?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
2
82
How are we going to verify authenticity?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
2
83
Who is going to verify the classification?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
2
84
Where can people manually verify if it is AI generated?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
2
85
What are the implications beyond the exact content? e.g job security for creatives?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
1
86
What types of risks can arise from transparency? (think privacy issues)
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
1
87
What does “authentic” content / information mean?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
88
Who might care the most about transparency? Users or providers and why?
Authenticity, Provenance, and Transparency
Trust, Authenticity, and User Empowerment
89
How will different users interpret signals differently?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
1
90
Where should users go if they have a complaint?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
91
How does disclosure impact people’s trust and perceived credibility?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
92
How to responsibly disclose artistic uses of AI?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
93
Why disclosure is considered as key part of user satisfaction?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
94
What are the consequences of mis-disclosure?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
95
How much does the user care?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
96
Why would you choose this to communicate (any empirical studies to base your actions upon ?)
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
97
Why AI should be disclosed?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
98
Why would people care about AI disclosure if AI content is flawless or productive?
Building Trust in AI Systems
Trust, Authenticity, and User Empowerment
99
How can we train users to spot authentic content without relying on explicit cues?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
2
100
What type of information empowers/triggers “act of resistance”?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
2
101
Why can’t people distinguish themselves?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
2
102
Where can people report false classification?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
1
103
Where should we get ask if we suspect things (images, texts) are generated by AI?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
104
How to balance AI generated “True” and “fake”?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
105
What could be the most obvious feature for fake/AI generated content?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
106
What are the features for fake/generated content?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
107
What is the tolerated degree of user satisfaction for generated content?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
108
Who could benefit from AI Disclosure?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
109
Why people would care about AI disclosure or not?
Empowering Users through Education & Awareness
Trust, Authenticity, and User Empowerment
110
Why do users need to know they are interacting with AI?
User Responsibility and Agency
Trust, Authenticity, and User Empowerment
3
111
What information should the users acquire? Are they responsible for it?
User Responsibility and Agency
Trust, Authenticity, and User Empowerment
1
112
Why is it important for promoting democracy?
User Responsibility and Agency
Trust, Authenticity, and User Empowerment
113
Who is going to continue using AI to create Fake content?
User Responsibility and Agency
Trust, Authenticity, and User Empowerment
114
When to fully avoid the content without AI disclosure?
User Responsibility and Agency
Trust, Authenticity, and User Empowerment
115
How much freedom should the users get to (not) see the classification?
Personalization and User Preferences
User Experience, Information Overload, and Personalization
3
116
How do users prefer to be informed?
Personalization and User Preferences
User Experience, Information Overload, and Personalization
2
117
How to adapt AI disclosures personally?
Personalization and User Preferences
User Experience, Information Overload, and Personalization
1
118
How to present them in different devices?
Personalization and User Preferences
User Experience, Information Overload, and Personalization
119
How to allow users to change the personalized UI settings?
Personalization and User Preferences
User Experience, Information Overload, and Personalization
120
Why do users trust AI more/less than a human?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
3
121
How would the disclosure affect the users?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
1
122
What is the psychological difference between AI and Human?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
123
What are the differences users feel with fake or deep fake content?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
124
Who are sensitive or insensitive for AI disclosure?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
125
What is the criteria for the provider for balancing AI content and human content?
Psychology of Human-AI Interaction
User Experience, Information Overload, and Personalization
126
What is a reliable method to test the classification?
Standardization
User Experience, Information Overload, and Personalization
3
127
Where on the website should Al be disclosed? Terms of service, every piece of content, every element?
Standardization
User Experience, Information Overload, and Personalization
1
128
Who determines how disclosure should be provided, according to which standards?
Standardization
User Experience, Information Overload, and Personalization
1
129
What would ensure that users are aware of the disclosure? (How will it be made? Another extensive terms and conditions or cookies? Do these currently work?
Standardization
User Experience, Information Overload, and Personalization
130
Who should assess if the AI disclosure is real/true?
Standardization
User Experience, Information Overload, and Personalization
131
How do you decide that the certification was designed by the right people and represents all relevant criteria?
Standardization
User Experience, Information Overload, and Personalization
132
Who communicates how the interaction and what stage of the interaction does this take place?
Standardization
User Experience, Information Overload, and Personalization
133
Where is the classification going to be displayed?
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
4
134
Where is the balance between communicating enough and information overload?
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
2
135
How can you effectively communicate use of Al without distracting from content?
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
1
136
How can we make trust labels meaningful?
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
1
137
How to balance/control the information overload?
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
138
What would the effect of disclosure be? e.g., a page full of AI warnings may be off putting for the user
User Interfaces and Information Overload
User Experience, Information Overload, and Personalization
139
Who are you communicating to? e.g., different people may require different types of communication mechanisms
User-Centric Information Design
User Experience, Information Overload, and Personalization
4
140
What under information is meaningful for the recipient?
User-Centric Information Design
User Experience, Information Overload, and Personalization
2
141
What format the users would like to see about the disclosure?
User-Centric Information Design
User Experience, Information Overload, and Personalization
2
142
Where to provide the AI disclosure? Before the generated content or after that?
User-Centric Information Design
User Experience, Information Overload, and Personalization
2
143
Where in people’s living space should nature of whose clearly be communicated? (e.g, policy notice, screen)?
User-Centric Information Design
User Experience, Information Overload, and Personalization
1
144
How do you decide what type of information to communicate and who is your target group ?
User-Centric Information Design
User Experience, Information Overload, and Personalization
1
145
What is the balance between AI generated “fake” and “true”?
User-Centric Information Design
User Experience, Information Overload, and Personalization
146
How to disclose it (AI/human made) by an implicit tag?
User-Centric Information Design
User Experience, Information Overload, and Personalization
147
How can we evaluate if the disclosure is sufficient enough information for the users?
User-Centric Information Design
User Experience, Information Overload, and Personalization
148
When do you engage in interdisciplinary collaborations between lawyers (so that you what and when to communicate) and those designing communications (so that lawyers know what’s possible and propose policy changes)
User-Centric Information Design
User Experience, Information Overload, and Personalization
149
What kind of user feedback/ input is incorporated into the AI system as disclosed?
User-Centric Information Design
User Experience, Information Overload, and Personalization
[…]

### 13 Mar 2024 v2

@@
 We believe these can not only help inform future legal developments and interpretations of Article 52, but also provide a starting point for Human-Computer Interaction research to (re-)examine disclosure transparency from a human-centered AI lens.
-EU AI Act, Article 52, generative artificial intelligence, disclosures, transparency, questions
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)
+EU AI Act, Article 52, generative artificial intelligence, disclosures, transparency, obligations, law, research questions
+journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613905.3650750isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing Human computer interaction (HCI)
 1. Introduction
-“Fake realities will create fake humans. Or, fake humans will generate fake realities
-and then sell them to other humans, turning them, eventually, into forgeries of themselves.”
+“Fake realities will create fake humans. Or, fake humans will generate fake realities and then sell them to other humans, turning them, eventually, into forgeries of themselves.”
 — Philip K. Dick, I Hope I Shall Arrive Soon, 1980
@@
 Foundation models, developed on the basis of deep neural networks and self-supervised learning, have gained widespread acceptance (Bommasani et al., 2021).
-These models have led to the emergence of Generative AI (GenAI) tools like Midjourney and ChatGPT, and have shown impressive proficiency in producing media content, including images (Saharia et al., 2022), text (OpenAI, 2023), videos (Ho et al., 2022), and audio (Borsos et al., 2022).
+These models have led to the emergence of Generative AI (GenAI) tools like Midjourney and ChatGPT, and have shown impressive capabilities in producing media content, including images (Saharia et al., 2022), text (OpenAI, 2023), videos (Ho et al., 2022), and audio (Borsos et al., 2022).
 These have the potential to drastically impact users and the media sector, essentially blurring the line between fiction and reality as users engage with media.
@@
 In fact, in the European Union (EU), such an obligation will be incorporated into the upcoming AI Act (Act, 2022).
-Despite that the AI Act is still in development and subject to change, the core obligation to “disclose AI-driven interaction” appears to be a stable one.
+Despite that the AI Act is still in development and subject to change, the core obligation to “disclose AI-driven interaction” appears to be stable.
 Algorithmic approaches to automatically detect text (co-)produced using GenAI tools (e.g., ChatGPT) are so far unreliable (Sadasivan et al., 2023), and can result in consistent biases against specific user groups (e.g., non-native English writers (Liang et al., 2023)).
@@
 Despite concerted efforts worldwide, constructing even simple disclosure measures (e.g., watermarks) that actually succeed in reducing public risks is itself a difficult task.
-Disclosure is defined as “to make known or public” (Webster, 2024), where in the context of AI, disclosures can “contain information about the data collection, data processing, and decision-making practices of a digital product and are voluntarily provided by the product’s vendor (an individual developer or an organization).” (Open Ethics Initiative (Lukianets, [n.d.])).
+Disclosure is defined as “to make known or public” (Webster, 2024), where in the context of AI, disclosures can “contain information about the data collection, data processing, and decision-making practices of a digital product and are voluntarily provided by the product’s vendor (an individual developer or an organization).” (Open Ethics Initiative (Lukianets, 2024)).
 A fundamental challenge here is in ensuring effective transparency measures, given the rapid pace of AI development and deployment.
-To mitigate harms and risks, regulatory efforts such as the AI Disclosure Act of 2023 in the United States (Torres, 2023) and the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) aim to tackle this.
+To mitigate harms and risks, regulatory efforts such as the AI Disclosure Act of 2023 in the United States (Torres, 2023) and the EU AI Act (Act, 2022) (currently under discussion (of the EU, 2024)) aim to tackle this.
 In this work, we focus on the EU AI Act proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses the issue of AI system transparency [^2] Even though the final text of the AI Act must still be agreed upon, the inclusion of a ”disclosure obligation” is (almost) a certitude..
@@
 2.1. EU AI Act’s Article 52 and AI disclosure implications
-We focus on the EU AI Act (Act, 2022) (currently under discussion (of the EU, [n.d.])) proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses AI system transparency.
+We focus on the EU AI Act (Act, 2022) (currently under discussion (of the EU, 2024)) proposition (Article 52: “Transparency obligations for providers and users of certain AI systems”), that addresses AI system transparency.
 It raises two important clauses:
@@
 For example, ‘what’, ‘how’, or ‘when’ should such AI disclosures take place, even if automatically detected, remains unclear.
-While the most recent draft of the AI Act (Monitor, [n.d.]) touches on aspects of this, where Art.
+While the most recent draft of the AI Act (Monitor, 2024) touches on aspects of this, where Art.
 52(3) (second subparagraph) highlights when to disclose, much remains uncharted regarding human oversight processes when exceptions are made.
@@
 Article 52 and Article 3 were printed on an A4 landscape paper, and placed near each flip chart for quick reference.
-Article 3 provides definitions of ‘provider’ and ‘user’ [^3] In the latest draft version of the EU AI Act (leaked online on January 21, 2024) (Monitor, [n.d.]), ‘user’ is now referred to as ‘deployer’.
+Article 3 provides definitions of ‘provider’ and ‘user’ [^3] In the latest draft version of the EU AI Act (leaked online on January 21, 2024) (Monitor, 2024), ‘user’ is now referred to as ‘deployer’.
 However, this does not impact our workshop outcomes, even if participants interpreted ‘user’ as intended by the law (‘deployer‘) or as ‘recipient’., that helped participants have a clear understanding of what these ‘legal’ terms mean.
@@
 Participants were offered candy in the session.
-At the end of each session, participants were thanked for their time and efforts.
+When the session ended, participants were thanked for their time and efforts.
 3.4. Participants
@@
 The sub-themes here touch upon important prior work that aims to address authenticity and trust, in the context of media.
+Liao & Sundar (Liao and Sundar, 2022) introduced a conceptual model called MATCH that describes how trustworthiness is communicated in AI systems through trustworthiness cues, and propose a requirement checklist to support technology creators in identifying reliable cues.
 Scharowski et al.
@@
 5.3. Conclusion
-Our work contributes 149 questions clustered into five themes and 18 sub-themes, that we believe can assist us in tackling the challenge of transparently communicating AI disclosures – for media providers, as well as end-users consuming media content everyday across platforms and devices.
-We hope our questions underscore the importance for better understanding user needs and reactions to transparency obligations, and support establishing user-centric designs for transparent AI disclosures that ultimately foster democratic societies based on truth rather than AI-generated fiction.
+Our work contributes 149 questions clustered into five themes and 18 sub-themes, that we believe can assist in tackling the challenge of transparently communicating AI disclosures – for media providers, as well as end-users consuming media content everyday across platforms and devices.
+We hope our questions underscore the importance of better understanding user needs and reactions to transparency obligations, and support establishing user-centric designs for transparent AI disclosures that ultimately foster democratic societies based on truth rather than AI-generated fiction.
 Acknowledgements.